OpenStudy (anonymous):

i would like help 410 divided by 41+40+36-14

OpenStudy (tylerd):

\[\frac{ 410 }{ 41+40+36-14 }\]

OpenStudy (tylerd):

add the ones on the bottom first

OpenStudy (tylerd):

41+40+36-14=?

OpenStudy (anonymous):

i got 103

OpenStudy (tylerd):

so 410/103=?

OpenStudy (anonymous):

yep and i got .2512

OpenStudy (tylerd):

um its around 3.98

OpenStudy (tylerd):

if ure using ure calc u do 410 first then hit the divide button second.